OpenStudy (anonymous):

Which of the following is equivalent to (1/4+y)^4? a) (16+8y+y^2)(y^2+8y+16) b) (y^3+1/4y^2+1/16y+1/64)(1/4+y) c) (y^3+3/4y^2+3/16y+1/64)(1/4+y) d) (4+y)^2(4+y)^2

OpenStudy (anonymous):

what do you think? :)

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

\[(\frac{1}{4}+y)^4=(\frac{1}{4}+y)^3(\frac{1}{4}+y)\]expand the \((\frac{1}{4}+y)^3\) part using formula\[(a+b)^3=a^3+3a^2b+3ab^2+b^3\]and see what happens :)

OpenStudy (anonymous):

no clue lol im not good at this

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thanx